OpenStudy (i_love_my_nieces):

The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show all of your work.

OpenStudy (whpalmer4):

What is the formula for the volume of a cylinder? What is the formula for the surface area of a cylinder?

OpenStudy (i_love_my_nieces):

Dude stop following me!!!!

OpenStudy (whpalmer4):

The purpose of OpenStudy is NOT to give you the answers, it's to show you how to find them yourself.

OpenStudy (i_love_my_nieces):

Okay, but stp following me!!

OpenStudy (whpalmer4):

Go back to the previous post if you have already forgotten what you were told, get the formulas, plug in the numbers. That's all there is to it.

OpenStudy (anonymous):

surface of a cylinder: \[A = 2*\pi*r^2+2*\pi*r*h\] Where r is radius and is height Volume: \[V = \pi*r^2*h\] Plug in numbers and you're done :)